# refs.bib {refsection}

### v2

@@ Abstract
With the growing use of language models (LMs) in clinical environments, there is an immediate need to evaluate the accuracy and safety of LM-generated medical text. Currently, such evaluation relies solely on manual physician review. However, detecting errors in LM-generated text is challenging because 1) manual review is costly and 2) expert-composed reference outputs are often unavailable in real-world settings. While the “LM-as-judge” paradigm (a LM evaluating another LM) offers scalable evaluation, even frontier LMs can miss subtle but clinically significant errors. To address these challenges, we propose MedVAL, a self-supervised framework that leverages synthetic data to train evaluator LMs to assess whether LM-generated medical outputs are factually consistent with inputs, without requiring physician labels or reference outputs. To evaluate LM performance, we introduce MedVAL-Bench, a dataset of 840 physician-annotated outputs across 6 diverse medical tasks capturing real-world challenges, including a multilingual task reviewed by bilingual physicians. Each output is reviewed following a physician-defined taxonomy of risk levels and error categories, enabling evaluation of LMs in making safety decisions for deployment. Across 10 state-of-the-art LMs spanning open-source, proprietary, and medically adapted models, MedVAL fine-tuning significantly improves ($p<0.001$) alignment with physicians on both seen and unseen tasks, increasing average F1 scores from 66% to 83%, with per-sample safety classification scores up to 86%. MedVAL improves the performance of even the best-performing proprietary LM (GPT-4o) by 8%. To support a scalable, risk-aware pathway towards clinical integration, we open-source the 1) codebase, 2) MedVAL-Bench, and 3) MedVAL-4B, the best-performing open-source LM. Our research provides the first evidence of LMs approaching expert-level validation ability for medical text.
@@ 1 Introduction
Language models (LMs) are increasingly leveraged for documentation in medical settings, supporting tasks such as clinical text summarization, report generation, or question answering (golob2016painful; moradi2022deep; singhal2023large; van2024adapted; aali2025dataset; tanno2025advancing; singhal2025toward). While LMs offer potential for reducing documentation burden sinsky2016allocation; gesner2019burden; ratwani2018usability; ehrenfeld2018technology, they are not infallible and frequently generate plausible text that contains subtle errors such as hallucinations, omissions, or certainty misalignments pal2023med; chang2025red; luo2024factual. This issue is critical within medicine, where errors are often cloaked in jargon, misleading even experienced practitioners kim2025medical.
@@
The adoption of LMs brown2020language; zhao2023survey; bubeck2023sparks for medical applications necessitates reliable risk assessment through medical text validation, which involves determining whether an output from a LM is factually consistent with the input. Currently, this process requires manual physician review both pre- and post-deployment to ensure continued safety, contributing to documentation workload and associated cognitive fatigue for physicians yackel2010unintended; gershanik2011critical; bowman2013impact; sasseville2025impacts.
@@
Despite this need, automatically evaluating LM-generated outputs remains difficult to scale. Traditional NLP metrics fail to capture subtle, clinically significant errors that can impact patient care banerjee2024rexamine. These metrics rely on expert-composed outputs (reference outputs), which are not available in many real-world scenarios zhou2023survey. Moreover, manual physician review remains the reference standard for validating medical text, but it is costly, time-consuming, and not scalable. With physician shortages and burnout at an all-time high arndt2017tethered; shanafelt2016relationship; robinson2018novel; toussaint2020design, assigning physicians additional validation tasks is impractical. Therefore, automated and scalable strategies that minimize reliance on expert-driven data are urgently needed to adapt LMs for medical text validation.
@@
Using a LM to evaluate another LM’s output, also termed the "LM-as-judge" paradigm, has shown promise in automating the evaluation of LM-generated text. However, most general-purpose LM-as-judge methods treat LMs as static evaluators gu2024survey, often lacking the granularity required to assess nuanced, high-stakes medical text min2023factscore; zha2023alignscore; wang2024calibrating; band2024linguistic; liu2024enhancing. Recent medical-specific LM-as-judge approaches suffer from certain practical limitations; these methods often assume resources like expert-labeled training data mehenni2025medhal, availability of reference outputs xie2023doclens, or retrieval-based evidence chung2025verifact, none of which are readily scalable. On the other hand, many approaches focus on subdomains such as radiology (e.g., chest X-rays), limiting generalization to broader medical tasks hardy2024rextrust; rao2024rexerr; ostmeier2024green; huang2024fineradscore; zambrano2025clinically; yu2023evaluating.
@@
To overcome these challenges, we introduce MedVAL (medical text validator), a self-supervised framework (Figure 1) that leverages synthetic data to train LMs for robust medical text validation. Our framework involves curating high-quality synthetic training examples by leveraging the agreement between a generator and a validator LM as a proxy for physician judgment li2023benchmarking. Through this training process, MedVAL enables LMs to assess whether an output is factually consistent with the input by assigning one of four risk levels, while flagging "unsafe for deployment" outputs at near physician-level reliability. In comparison to related methods, our framework enables: 1) development of scalable evaluators without physician-in-loop supervision, 2) evaluation without reliance on reference outputs, and 3) adaptation across diverse medical tasks.
@@
We evaluate MedVAL across 6 open-source and 4 proprietary LMs on MedVAL-Bench, our newly created dataset consisting of 840 physician-annotated outputs, where each output is annotated following a physician-defined taxonomy of risk levels and error categories. MedVAL-Bench includes 6 diverse tasks (focused on LM output evaluation), including medication question answering, patient query summarization, radiology report summarization, radiology impression simplification, hospital course translation, and doctor-patient dialogue summarization. To reflect the multilingual nature of U.S. clinical practice, where Spanish is the second most spoken language amn2023languages, we include an English-to-Spanish translation task annotated by bilingual physicians.
@@
Unlike prior LM-as-judge methods that rely on zero-shot inference, we demonstrate that MedVAL fine-tuning improves the validation capabilities of all underlying LMs across in-distribution and out-of-distribution tasks. MedVAL yields significant gains ($p<0.001$): average baseline F1 scores for four-class risk-level classification improve from 36.7% to 51.0%, and for binary safe/unsafe classification, from 66.2% to 82.8%. Overall, MedVAL marks a significant step toward closing the gap between automated and expert-level assessment.
@@ 2.1 MedVAL Training
We propose a self-supervised framework for training LMs to validate LM-generated medical text, involving three stages: 1) synthetic data generation, 2) data filtering, and 3) fine-tuning. Figure 2 describes the data curation process, and Supplementary Algorithm 1 summarizes the complete training process.
@@
We denote the generator LM as $g_{\theta}$, validator LM as $v_{\phi}$, the fine-tunable validator LM as $v_{\alpha}$, the input as $x$, the corresponding reference output as $y$, the output as $\hat{y}=g_{\theta}(x)$ or $\hat{y}=y$ (if $y$ available), and the perturbed output as $\hat{y}_{\delta}=g_{\theta}(x,\delta)$. Here, $\delta\in[0,1]$ represents the perturbation level in terms of factual degradation, where $\delta=0$ represents an unperturbed generation and higher $\delta$ values correspond to more factually inconsistent generations. $\hat{\delta}_{clean}=v_{\phi}(x,\hat{y})$ and $\hat{\delta}_{corrupt}=v_{\phi}(x,\hat{y}_{\delta})$ denote the validator’s prediciton of the factual degradation in the clean $\hat{y}$ and perturbed $\hat{y}_{\delta}$ outputs, respectively.
@@
$\delta=0\%$The output should contain no clinically meaningful factual inconsistencies. Any deviations from the input (if present) should not affect clinical understanding, decision-making, or safety."
$\delta=33\%$The output should contain subtle or ambiguous inconsistencies that are unlikely to influence clinical decisions or understanding. These inconsistencies should not introduce confusion or risk."
$\delta=67\%$The output should contain inconsistencies that could plausibly affect clinical interpretation, documentation, or decision-making. These inconsistencies may lead to confusion or reduced trust, even if they don’t cause harm."
$\delta=100\%$The output should include one or more inconsistencies that could result in incorrect or unsafe clinical decisions. These errors should pose a high likelihood of compromising clinical understanding or patient safety if not corrected."
@@
The first term in $\mathcal{M}_{\text{absolute}}$ rewards clean generations that achieve validator scores close to 0, and the second term rewards perturbed generations whose validator scores closely match the expected degradation $\delta$. $\mathcal{M}_{\text{relative}}$ ensures that the difference between the predicted scores reflects the expected degradation. To bound the score between 0 and 1 ($\uparrow$ score $=$ $\uparrow$ generator-validator consistency), we divide it by its maximum possible value:
@@
We create our synthetic dataset $\mathcal{D}_{\text{train}}=\{x,\hat{y},\hat{\delta}_{\text{clean}},\hat{y}_{\delta},\hat{\delta}_{\text{corrupt}}\}$ by filtering examples from $\mathcal{D}$ where $\mathcal{M}_{\text{MedVAL}}\geq\tau$ (pre-specified threshold). We choose $\tau$ to retain only examples with high generator-validator agreement. We then fine-tune an arbitrary LM $v^{*}_{\alpha}=\textit{SFT}(v_{\alpha},\mathcal{D}_{\text{train}})$ using a standard parameter-efficient supervised fine-tuning (SFT) objective. To improve robustness, we optionally ensemble multiple MedVAL fine-tuned LMs by aggregating their predictions and applying post-hoc thresholding.
@@
Task Level 1 (%) Level 2 (%) Level 3 (%) Level 4 (%) N
medication2answer 36.3% 11.1% 22.2% 30.4% 135
query2question 38.3% 20.8% 28.3% 12.5% 120
report2impression 34.2% 19.5% 21.1% 25.3% 190
impression2simplified 32.6% 30.0% 14.7% 22.6% 190
bhc2spanish 38.3% 11.7% 9.2% 40.8% 120
dialogue2note 47.1% 12.9% 15.3% 24.7% 85
Overall 36.7% 18.9% 18.6% 25.8% 840
@@ 2.2 Data
We introduce MedVAL-Bench (Table 2), a dataset for evaluating the ability of automated tools to assess factual consistency of LM-generated outputs (not their ability to generate input $\to$ output). The train set is self-supervised, containing inputs used to generate synthetic supervision labels, and reference outputs if available (medication2answer, query2question, report2impression, and dialogue2note). To enable LM benchmarking, the test set contains physician assessments of outputs spanning broad error severities. Next, we describe: 1) train/test input sources for each task, and 2) test-time data curation via the reader study.
@@ 2.2.1 Input Sources
The medication2answer task involves answering a medication-related question, and utilizes the MedicationQA dataset medicationqa, which contains real-world consumer health questions about medications and expert-written answers. Unlike other tasks, the input (a standalone question) often lacks sufficient context to validate the output’s factual consistency. Hence, we aim to evaluate whether LMs can reliably perform safety-critical decisions using their knowledge base and the original question. Due to the unavailability of pre-specified train/test splits, we randomly create the splits as specified in Table 2.
@@
The query2question task involves creating a concise query that encapsulates the essential details necessary to address the original question. We utilize the MeQSum dataset MeQSum, which comprises: 1) health-related queries from patients, sourced from messages sent to the U.S. National Library of Medicine, and 2) corresponding abbreviated questions formulated by three medical experts to ensure that the summaries facilitate complete and accurate answer retrieval. For our study, we sample a subset of examples from both the train and test splits defined by van2024adapted, as detailed in Table 2.
@@
The report2impression task involves processing the "findings" section of a radiology exam report to create an "impression" section that succinctly highlights the critical information. We leverage the Open-i dataset demner2016preparing, which consists of de-identified narrative chest x-ray reports from the Indiana Network for Patient Care database. We randomly sample train/test examples following the splits identified by van2024adapted, ensuring the number of train/test samples indicated in Table 2.
@@
The impression2simplified task involves simplifying radiology report impression sections to produce patient-friendly rewrites that preserve the clinical meaning while being understandable to a layperson. For this task, we extract the "impression" sections from the MIMIC-IV dataset johnson2020mimic, which contains comprehensive radiology reports sourced from patient stays in critical care units of the Beth Israel Deaconess Medical Center. Our sampling process ensures an equal distribution of imaging modalities that reflect the top imaging indications at Stanford Hospital, as confirmed by consultations with hospitalists. These include chest x-rays, CT scans of the abdomen and pelvis, CT scans of the head, MR brain studies, ultrasound exams of the pelvis, digital screening mammography, and transabdominal/transvaginal pelvic ultrasounds. Testing examples were selected via stratified random sampling as indicated in Table 2. Because MIMIC-IV johnson2020mimic is partially open-source, and to preserve this task for out-of-distribution evaluation, we select training examples from Open-i radiology report findings on a related but distinct report2simplified task (i.e., simplifying the findings section). For this task, we randomly sample training examples from the training split by van2024adapted, ensuring no overlap with the samples used in report2impression.
@@
The bhc2spanish task involves translating brief hospital course (BHC) sections from discharge summaries from English into Spanish. We employ the MIMIC-IV-BHC dataset aali2024mimic, which contains curated BHC sections from discharge summaries, written by healthcare providers at the Beth Israel Deaconess Medical Center. The MIMIC-IV-BHC dataset includes diverse patient encounters, ensuring a comprehensive representation. We randomly sample testing examples as indicated in Table 2.
@@
The objective of dialogue2note is to summarize doctor-patient conversations into an "assessment and plan". We utilize the ACI-Bench dataset yim2023aci; abacha2023overview; MEDIQA-Sum2023, which provides a collection of 1) 207 doctor-patient conversations and 2) patient visit notes. To ensure consistency, we randomly sample the testing examples (indicated in Table 2), from the test set defined by van2024adapted.
@@ 2.2.2 Physician Reader Study
We design our reader study to enable a robust evaluation of how well LMs assess the factual consistency of outputs, compared to physicians. For each test-time input, an LM-generated output was produced by sampling a perturbation level $\delta\in[{0,0.33,0.67,1.0}]$ uniformly at random, and injecting the corresponding perturbation instruction into the prompt as described in Table 1. This ensured that the test set contained a balanced mix of outputs with no risk, low risk, moderate risk, and high risk errors. Since our objective was to evaluate the validator’s ability to distinguish clinically meaningful degradation, we do not explicitly prompt the generator to produce only no-risk outputs, as doing so would instead assess generation quality. This design was necessary to allow evaluation of a validator’s ability to distinguish degradation across a meaningful severity spectrum. Importantly, these perturbations apply only to the test set, as the synthetic training set is curated using the pipeline in Figure 2. The prompts used for output generation are shown in Table 2.

### v4

# Toward expert-level medical text validation with language models

@@ Abstract
With the growing use of language models (LMs) in clinical environments, there is an immediate need to evaluate the accuracy and safety of LM-generated medical text. Currently, such evaluation relies solely on manual physician review. However, detecting errors in LM-generated text is challenging because 1) manual review is costly and 2) expert-composed reference outputs are often unavailable in real-world settings. While the “LM-as-judge” paradigm (a LM evaluating another LM) offers scalable evaluation, even frontier LMs can miss subtle but clinically significant errors. To address these challenges, we propose MedVAL, a novel, self-supervised, data-efficient distillation method that leverages synthetic data to train evaluator LMs to assess whether LM-generated medical outputs are factually consistent with inputs, without requiring physician labels or reference outputs. To evaluate LM performance, we introduce MedVAL-Bench, a dataset of 840 physician-annotated outputs across 6 diverse medical tasks capturing real-world challenges, including a multilingual task reviewed by bilingual physicians. Each output is reviewed following a physician-defined taxonomy of risk levels and error categories, enabling evaluation of LMs in making deployment safety decisions. Across 10 state-of-the-art LMs spanning open-source, proprietary, and medically adapted models, MedVAL distillation significantly improves ($p<0.001$) alignment with physicians across seen and unseen tasks, increasing average F1 scores from 66% to 83%. Despite strong baseline performance, MedVAL improves the best-performing proprietary LM (GPT-4o) by 8% without training on physician-labeled data, demonstrating a performance statistically non-inferior to a single human expert ($p<0.001$). To support a scalable, risk-aware pathway towards clinical integration, we open-source: 1) Codebase, 2) MedVAL-Bench, 3) MedVAL-4B. Our benchmark provides evidence of LMs approaching expert-level ability in validating AI-generated medical text.
@@ 1 Introduction
Language models (LMs) are increasingly leveraged for documentation in medical settings, supporting tasks such as clinical text summarization, report generation, or question answering [1, 2, 3, 4, 5, 6, 7]. While LMs offer potential for reducing documentation burden [8, 9, 10, 11], they are not infallible and frequently generate plausible text that contains subtle errors such as hallucinations, omissions, or certainty misalignments [12, 13, 14]. This issue is critical within medicine, where errors are often cloaked in jargon, misleading even experienced practitioners [15].
@@
The adoption of LMs [16, 17, 18] for medical applications necessitates reliable risk assessment through medical text validation, which involves determining whether an output from a LM is factually consistent with the input. Currently, this process requires manual physician review both pre- and post-deployment to ensure continued safety, contributing to documentation workload and associated cognitive fatigue for physicians [19, 20, 21, 22].
@@
Despite this need, automatically evaluating LM-generated outputs remains difficult to scale. Traditional NLP metrics fail to capture subtle, clinically significant errors that can impact patient care [23]. These metrics rely on expert-composed outputs (reference outputs), which are not available in many real-world scenarios [24]. Moreover, manual physician review remains the reference standard for validating medical text, but it is costly, time-consuming, and not scalable. With physician shortages and burnout at an all-time high [25, 26, 27, 28], assigning physicians additional validation tasks is impractical. Therefore, automated and scalable strategies that minimize reliance on expert-driven data are urgently needed to adapt LMs for medical text validation.
@@
Using a LM to evaluate another LM’s output, also termed the "LM-as-judge" paradigm, has shown promise in automating the evaluation of LM-generated text. However, most general-purpose LM-as-judge methods treat LMs as static evaluators [29], often lacking the granularity required to assess nuanced, high-stakes medical text [30, 31, 32, 33, 34]. Recent medical-specific LM-as-judge approaches suffer from certain practical limitations; these methods often assume resources like expert-labeled training data [35], availability of reference outputs [36], or retrieval-based evidence [37], none of which are readily scalable. On the other hand, many approaches focus on subdomains such as radiology (e.g., chest X-rays), limiting generalization to broader medical tasks [38, 39, 40, 41, 42, 43, 44].
@@
To overcome these challenges, we introduce MedVAL (medical text validator), a self-supervised, data-efficient distillation method (Figure 1), which curates high-quality synthetic examples by leveraging the agreement between a generator and a validator LM as a proxy for physician judgment [45]. Through this training, MedVAL enables LMs to assess whether an output is factually consistent with the input by assigning one of four risk levels, while flagging "deployment unsafe" outputs at near physician-level reliability. In comparison to prior methods, our framework enables: 1) development of scalable evaluators without physician-in-loop supervision, 2) evaluation without reliance on reference outputs, and 3) adaptation across diverse tasks.
@@
We evaluate MedVAL across 6 open-source and 4 proprietary LMs on MedVAL-Bench, our newly created dataset consisting of 840 physician-annotated outputs, where each output is annotated following a physician-defined taxonomy of risk levels and error categories. MedVAL-Bench includes 6 diverse tasks (focused on LM output evaluation), including medication question answering, patient query summarization, radiology report summarization, radiology impression simplification, hospital course translation, and doctor-patient dialogue summarization. To reflect the multilingual nature of U.S. clinical practice, where Spanish is the second most spoken language [46], we include an English-to-Spanish translation task annotated by bilingual physicians.
@@
Unlike prior LM-as-judge methods that rely on zero-shot inference, MedVAL distillation improves the validation capabilities of all underlying LMs across seen and unseen tasks. MedVAL yields significant gains ($p<0.001$): average baseline F1 scores for four-class risk-grading improve from 36.7% to 51.0%, and for binary safe/unsafe classification, from 66.2% to 82.8%. Our benchmark provides evidence that LMs can achieve performance statistically non-inferior to a single human expert on risk-based validation decisions ($p<0.001$).
@@ 2.1 MedVAL Training
We propose a self-supervised distillation method for training LMs to validate LM-generated medical text, involving three stages: 1) synthetic data generation, 2) data filtering, and 3) fine-tuning. Figure 2 describes the data curation process, and Supplementary Algorithm 1 summarizes the complete training process.
@@
[…]
@@
$\delta=0\%$ "The output should contain no clinically meaningful factual inconsistencies. Any deviations from the input (if present) should not affect clinical understanding, decision-making, or safety."
$\delta=33\%$ "The output should contain subtle or ambiguous inconsistencies that are unlikely to influence clinical decisions or understanding. These inconsistencies should not introduce confusion or risk."
$\delta=67\%$ "The output should contain inconsistencies that could plausibly affect clinical interpretation, documentation, or decision-making. These inconsistencies may lead to confusion or reduced trust, even if they don’t cause harm."
$\delta=100\%$ "The output should include one or more inconsistencies that could result in incorrect or unsafe clinical decisions. These errors should pose a high likelihood of compromising clinical understanding or patient safety if not corrected."
@@
The first term in $\mathcal{M}_{\text{absolute}}$ rewards clean generations that achieve validator scores close to 0, and the second term rewards perturbed generations whose validator scores closely match the expected degradation $\delta$. $\mathcal{M}_{\text{relative}}$ ensures that the difference between the predicted scores reflects the expected degradation. To bound the score between 0 and 1 ($\uparrow$ score == $\uparrow$ generator-validator consistency), we divide it by its maximum possible value:
@@
We create our synthetic dataset $\mathcal{D}_{\text{train}}=\{x,\hat{y},\hat{\delta}_{\text{clean}},\hat{y}_{\delta},\hat{\delta}_{\text{corrupt}}\}$ by filtering examples from $\mathcal{D}$ where $\mathcal{M}_{\text{MedVAL}}\geq\tau$ (pre-specified threshold). Importantly, we apply a single-pass consistency filter (not iterative), and choose $\tau$ to retain examples with high generator-validator agreement. We then fine-tune an arbitrary LM $v^{*}_{\alpha}=\textit{SFT}(v_{\alpha},\mathcal{D}_{\text{train}})$ using standard parameter-efficient supervised fine-tuning (SFT). To improve robustness, we optionally ensemble multiple MedVAL fine-tuned LMs by aggregating their predictions.
@@ 2.2 Data
We introduce MedVAL-Bench (Table 2), a dataset for evaluating the ability of LMs to validate LM-generated medical text. The train set is self-supervised, containing inputs (used to generate supervision labels) and reference outputs (if available). The test set additionally contains gold-standard physician assessments of outputs. Next, we describe: 1) train/test input sources, and 2) test-time data curation via the reader study.
@@ 2.2.1 Input Sources
The medication2answer task involves answering a medication-related question, and utilizes the MedicationQA dataset [47], which contains real-world consumer health questions about medications and expert-written answers. Unlike other tasks, the input (a standalone question) often lacks sufficient context to validate the output’s factual consistency. Hence, we aim to evaluate whether LMs can reliably perform safety-critical decisions using their knowledge base and the original question. Due to the unavailability of pre-specified train/test splits, we randomly create the splits as specified in Table 2.
@@
The query2question task involves creating a concise query that encapsulates the essential details necessary to address the original question. We utilize the MeQSum dataset [48], which comprises: 1) health-related queries from patients, sourced from messages sent to the U.S. National Library of Medicine, and 2) corresponding abbreviated questions formulated by three medical experts to ensure that the summaries facilitate complete and accurate answer retrieval. For our study, we sample a subset of examples from both the train and test splits defined by [4], as detailed in Table 2.
@@
The report2impression task involves processing the "findings" section of a radiology exam report to create an "impression" section that succinctly highlights the critical information. We leverage the Open-i dataset [49], which consists of de-identified narrative chest x-ray reports from the Indiana Network for Patient Care database. We randomly sample train/test examples following the splits identified by [4], ensuring the number of train/test samples indicated in Table 2.
@@
The impression2simplified task involves simplifying radiology report impression sections to produce patient-friendly rewrites that preserve the clinical meaning while being understandable to a layperson. For this task, we extract the "impression" sections from the MIMIC-IV dataset [50], which contains comprehensive radiology reports sourced from patient stays in critical care units of the Beth Israel Deaconess Medical Center. Our sampling process ensured an equal distribution of imaging modalities that reflect the top imaging indications at Stanford Hospital. These include chest x-rays, CT scans of the abdomen and pelvis, CT scans of the head, MR brain studies, ultrasound exams of the pelvis, digital screening mammography, and transabdominal/transvaginal pelvic ultrasounds. Test examples were selected via stratified random sampling as indicated in Table 2. Because MIMIC-IV [50] is partially open-source, we select training examples from Open-i radiology report findings on a related but distinct report2simplified task (i.e., simplifying the findings section). For this task, we randomly sample training examples from the training split by [4], ensuring no overlap with the samples used in report2impression.
@@
The bhc2spanish task involves translating brief hospital course (BHC) sections from discharge summaries from English into Spanish. We employ the MIMIC-IV-BHC dataset [51], which contains curated BHC sections from discharge summaries, written by healthcare providers at the Beth Israel Deaconess Medical Center. The MIMIC-IV-BHC dataset includes diverse patient encounters, ensuring a comprehensive representation. We randomly sample testing examples as indicated in Table 2.
@@
The objective of dialogue2note is to summarize doctor-patient conversations into an "assessment and plan". We utilize the ACI-Bench dataset [52, 53, 54], which provides a collection of 1) 207 doctor-patient conversations and 2) patient visit notes. To ensure consistency, we randomly sample the testing examples (indicated in Table 2), from the test set defined by [4].
@@ 2.2.2 Physician Reader Study
We design our reader study to enable a robust evaluation of how well LMs assess the factual consistency of outputs, compared to physicians. For each test-time input, an LM-generated output was produced by sampling a perturbation level $\delta\in[{0,0.33,0.67,1.0}]$ uniformly at random, and injecting the corresponding perturbation instruction into the prompt as described in Table 1. This ensured that the test set contained a balanced mix of outputs with no risk, low risk, moderate risk, and high risk errors. Since our objective was to evaluate the validator’s ability to distinguish clinically meaningful degradation, we do not explicitly prompt the generator to produce only no-risk outputs, as doing so would instead assess generation quality. This design was necessary to evaluate a validator’s ability to distinguish degradation across a meaningful severity spectrum. We observe that models sometimes diverge from requested perturbations; consequently, outputs often contain a blend of prompted and natural errors committed by LMs. In all cases, physician graders assign risk levels based solely on the observed text, independent of the prompt specification. Importantly, these perturbation rules apply to the test set, as the training set is curated using the pipeline in Figure 2.
@@
We present the error category distribution in Supplementary Figure 9. Across 840 examples, the most common error categories are fabricated claim (45.7%), missing claim (14.0%), and incorrect recommendation (12.6%). We also observe that error frequency rises with the risk grades (for grades 1–4, average 0.14 $\to$ 3.24 errors).